OpenStudy (anonymous):

dy/dx e^(cos(4t)+ln(t))

OpenStudy (anonymous):

dt/dx*

OpenStudy (anonymous):

\[ \left( -4\,\sin \left( 4\,t \right) +{t}^{-1} \right) {{\rm e}^{\cos \left( 4\,t \right) +\ln \left( t \right) }} \]